# A Survey of AI Scientists

### v4

@@ 3.2 Idea Generation
Idea generation transforms structured knowledge into testable hypotheses through scientific creativity, semantic grounding, and novelty control. We formalize this as a four-stage framework (Figure 5): conceptual fusion, hypothesis refinement, multi-agent brainstorming, and feasibility evaluation [Zhou2024HypothesisGenerationLLMs, su2024many].
@@
    Stage 1: Conceptual Fusion and Trend Extrapolation. Discovers novel research directions by identifying latent connections across disparate knowledge domains. Dense retrieval surfaces conceptually related papers, embedding-based clustering reveals thematic patterns, and knowledge graph reasoning [Zhou2024HypothesisGenerationLLMs] enables analogical transfer. Temporal citation analysis identifies emerging trends. This constructs novelty-annotated hypothesis banks with estimated originality scores.
@@
    Stage 2: Hypothesis Refinement and Knowledge-Grounded Pruning. Filters and refines candidate hypotheses through multi-criteria validation ensuring logical coherence and feasibility. Knowledge graph validation checks consistency with theoretical frameworks, provenance-aware distillation [vasu2025hyper] traces components to supporting evidence, and simulation-based testing evaluates tractability. Advanced systems employ reinforcement learning for iterative refinement and uncertainty-aware rejection sampling to eliminate speculative hypotheses.
@@
    Stage 3: Multi-Agent Brainstorming and Cross-Domain Recombination. Enhances hypothesis diversity through structured collaboration among specialized agents. Multi-role architectures assign distinct responsibilities—critic agents identify flaws, generator agents propose variations, verifier agents check consistency [su2024many, Ghafarollahi2024SciAgents]. Structured message passing enables agents to ground discussions in retrieved evidence while exploring complementary perspectives, producing hypothesis sets with conceptual diversity.
@@
    Stage 4: Feasibility and Novelty Evaluation. Ranks hypotheses through multi-dimensional scoring combining technical feasibility, scientific novelty, and potential impact. Graph-based semantic evaluation quantifies conceptual distance from existing work, while expert-annotated benchmarks [guo2025ideabench, liu2025researchbench] provide ground-truth assessments. High-scoring hypotheses receive priority routing to experimental preparation, establishing a quality-controlled pipeline from ideation to experimentation.
@@ 3.3 Experimental Preparation
    Stage 1: Experimental Framing. Translates hypotheses into actionable experimental plans by formalizing variables, metrics, and testable assumptions. Multimodal reasoning over tables and charts [wu2025tablebench] enables structured variable definition, while automated model template generation [Majumder2024DiscoveryBench] provides standardized experimental frameworks. This stage establishes clear objectives, measurable outcomes, and explicit constraints for subsequent execution.
@@
    Stage 3: Protocol Implementation and Plan Reasoning. Operationalizes experiments through mathematical modeling and multi-step execution planning. Intermediate checkpoints [guo2024ds] enable continuous introspection and error recovery, while dynamic schema adaptation allows real-time protocol adjustments. This stage transforms high-level experimental designs into executable sequences with built-in monitoring and adaptive re-execution capabilities.
@@
    Stage 4: Reproducibility and Lifecycle Tracking. Maintains comprehensive provenance documentation throughout experimental lifecycles. Parameter tracking, environmental logging [Gu2024BLADE], and versioning systems record all experimental configurations and execution details. This stage quantifies cross-agent consistency metrics and ensures verifiable scientific operations through cryptographic hashing, version control, and standardized metadata schemas.
@@ 3.4 Experimental Execution
Experimental execution operationalizes hypothesis-driven protocols into verifiable empirical outcomes through closed-loop reasoning and adaptive control [Boiko2023Coscientist, Curie2025, Kon2025EXPBench, Mandal2025AILA]. Unlike linear automation pipelines, this stage employs continuous feedback between symbolic planning and physical instrumentation, enabling real-time error correction and parameter optimization. We formalize this process through four sequential stages (Figure 7): Protocol Instantiation, Instrument Invocation, Adaptive Execution, and Data Validation.
@@
    Stage 1: Protocol Instantiation. Translates abstract experimental designs into executable, domain-specific procedures through language-to-action mapping. BioPlanner demonstrates natural-language-to-protocol conversion for biochemical validation [BioPlanner2023], while Curie formalizes modular representations enabling task decomposition [Curie2025]. Hierarchical frameworks encode safety constraints, device mappings, and procedural templates into structured abstraction layers [Shi2025HierEncapRepProtocol].
@@
    Stage 2: Instrument and Tool Invocation. Interfaces instantiated protocols with heterogeneous experimental infrastructure including robotic laboratories, simulators, and computational clusters. Self-driving laboratories [Tobias2025SDLReview] integrate robotic manipulation and closed-loop control, while multi-agent systems like ORGANA [Darvish2025ORGANA] and AutoLabs [Panapitiya2025AutoLabs] orchestrate apparatus scheduling. Autonomous agents demonstrate scalability through unified API integration and metadata-driven control.
@@
    Stage 3: Adaptive Execution and Feedback. Implements real-time monitoring and dynamic protocol revision through closed-loop control. EXP-Bench [Kon2025EXPBench] and Curie [Curie2025] leverage intermediate metrics to trigger parameter adjustments. DS-1000 [Lai2023DS1000] and notebook generation [Yin2022NL2Notebook] demonstrate code-testing feedback loops, while CoScientist [Boiko2023Coscientist] realizes physical closed-loop reasoning. These frameworks converge reinforcement learning, self-critique, and error detection into adaptive control.
@@
    Stage 4: Data Acquisition and Validation. Consolidates multimodal experimental outputs into provenance-linked, validated datasets through comprehensive metadata annotation and statistical verification [Mandal2025AILA]. Curie [Curie2025] and EXP-Bench [Kon2025EXPBench] embed runtime analytics validating outputs against simulations. Physics applications aggregate sensor telemetry for causal interpretation. This establishes an auditable evidence chain linking execution logs to scientific reporting.
@@ 3.5 Scientific Writing
Scientific writing transforms experimental findings into verifiable, publication-grade scholarly narratives through factual grounding, data–text alignment, and ethical compliance. LLM-assisted writing now permeates biomedical publishing, educational research [cheng2025artificial], and simulation science [cheng2025artificial]. We formalize this ecosystem through five interlinked stages (Figure 8): Drafting, Data–Text Linking, Peer-Review Automation, Ethical Governance, and Publication Optimization.
@@
    Stage 1: Drafting and Structural Composition. Converts analytical outputs into disciplinary-compliant academic prose through logical structuring and cross-sectional coherence [telenti2024large, khalifa2024using]. Domain-specific assistants achieve expert-level fluency via section-specific prompts [cheng2025artificial], while citation-aware mechanisms map claims to verified sources and structure-aware templates enforce parallelism [mugaanyi2024evaluation]. Human-AI co-editing workflows reduce grammatical noise and accelerate drafting [guo2024artificial].
@@
    Stage 2: Data–Text Linking and Multimodal Representation. Aligns quantitative data with narratives through automated cross-referencing of tables and figures [mugaanyi2024evaluation, WritingBench2025Benchmark]. Hybrid language-vision models generate visual elements from structured data [salvagno2023can], while grounded captioning mechanisms automate caption rewriting and figure suggestion [son2025ai]. This establishes scientific writing as multimodal composition where LLMs mediate between data, visuals, and text.
@@
    Stage 3: Peer-Review Automation and Quality Verification. Automates manuscript evaluation through LLM-based reviewers trained on editorial corpora, identifying inconsistent claims and logical fallacies [zhou2025large, son2025ai]. Systems emulate referee workflows via structured comments and quantitative scores, implementing iterative review loops that reduce latency [feyzollahi2025adoption]. Reproducibility and algorithmic bias remain challenges, necessitating transparent confidence calibration [telenti2024large].
@@
    Stage 4: Ethical Compliance and Authorship Governance. Enforces mandatory disclosure of AI involvement to prevent ghostwriting [ateriya2025exploring, resnik2025ethics, Mwita2025AIWritingEthics]. Standardized frameworks document prompt histories, model identifiers, and revision provenance [roe2023review], while layered attribution models recognize human intellectual contribution and algorithmic assistance [jo2023promise]. These mechanisms anchor trust and accountability in AI-augmented authorship.
@@
    Stage 5: End-to-End Publication Optimization. Integrates drafting, multimodal generation, review, and ethics into unified workflows. Frameworks like WritingBench [WritingBench2025Benchmark] and SPOT [son2025ai] evaluate manuscripts for factual accuracy and citation soundness. Adaptive agents monitor document lifecycles, suggesting figure adjustments, formatting references, and validating ethical disclosures, enabling scientists to transition to supervisory orchestration [khalifa2024using].
@@ 3.6 Paper Generation
Paper generation integrates ideation, experimentation, visualization, composition, and review into autonomous manuscript creation workflows (Figure 9). Systems like The AI Scientist v1/v2 generate complete papers—including figures, experiments, and reviews—with minimal human oversight [Lu2024AIScientist, Yamada2025AIScientistV2], while AI-Researcher establishes end-to-end pipelines and benchmarks (Scientist-Bench) for systematic evaluation [Tang2025AIResearcher]. We decompose this process into four stages: Manuscript Drafting, Visual & Tabular Composition, Review & Revision, and Publication Dissemination.
@@
    Stage 1: Manuscript Drafting. Transforms structured outputs into full manuscript sections through retrieval-augmented generation. The AI Scientist v1 converts experiment logs to LaTeX drafts, synthesizes literature context, and embeds citations automatically [Lu2024AIScientist]. AIGS extends this via multi-agent pipelines for full-cycle generation [Liu2024AIGS], employing template-based formatting for journal compliance.
@@
    Stage 2: Visual & Tabular Composition. Generates and integrates visual artifacts with narrative alignment. The AI Scientist converts experimental outputs to plots and tables, generating captions and inserting cross-references [Lu2024AIScientist], while SciSciGPT integrates data processing with figure generation [Shao2025SciSciGPT]. Core technologies include automatic plotting, figure-caption co-generation, and multimodal consistency validation.
@@
    Stage 3: Review & Revision Agent. Embeds internal reviewer agents to critique and refine manuscripts. The AI Scientist v2 implements agentic tree-search with dedicated reviewer components generating structured feedback [Yamada2025AIScientistV2], while AI-Researcher enables automated revision loops [Tang2025AIResearcher]. Key methodologies include peer-review-trained models, LLM-driven comment generation, and iterative author-reviewer simulation.
@@
    Stage 4: Publication Dissemination. Prepares manuscripts for submission through venue-specific formatting and metadata management. Automated pipelines ensure submission readiness [madanchian2024ai], while post-publication modules enable versioning and citation tracking. Core technologies include template-mapping tools, submission-API integration, metadata provenance logging, and ethical-disclosure verification.
@@ 4.1 General AI Scientist Systems
    The AI Scientist v1 [Lu2024AIScientist] deploys modular multi-agent architecture—planning, coding, analysis, writing—coordinated by a meta-scientist module, autonomously selecting topics, generating executable code, drafting papers, and reproducing canonical machine learning studies.
@@
    The AI Scientist v2 [Yamada2025AIScientistV2] advances v1 through agentic tree-search planning, dynamically exploring parallel hypotheses and evaluating novelty/validity via reflective feedback loops.
@@
    AI-Researcher [Tang2025AIResearcher] prioritizes transparency through provenance-tracking memory graphs recording all artifacts—code, data logs—co-developed with Scientist-Bench for reproducibility evaluation.
@@
    Curie [Curie2025] achieves rigorous experimental control via causality-aware planning loops, automating machine learning hypothesis testing with explicit causal assumptions.
@@ 4.2 Chemistry and Materials Science
    Coscientist [Boiko2023Coscientist] integrates GPT-4 reasoning with robotic liquid handlers, autonomously planning reactions, generating control code, interpreting spectroscopic feedback, and iteratively refining hypotheses.
@@
    A-Lab [Szymanski2023AutonomousLab] combines Bayesian optimization with LLM-guided experiment design, autonomously synthesizing and characterizing thousands of novel inorganic materials weekly.
@@
    Robotic AI Chemist [Song2024RoboticAIchemist] converges cognitive autonomy (literature-grounded reasoning) with physical autonomy (robotic manipulation) for end-to-end reaction design and execution.
@@
    AutoLabs [Panapitiya2025AutoLabs] deploys multi-agent self-correction cycles—planning, control, safety auditing—detecting anomalies and recalibrating instruments for enhanced throughput and safety.
@@ 4.3 Biology and Biomedical Research
    BioPlanner [BioPlanner2023] establishes benchmarks for LLM-driven biological protocol design, formalizing translation of research goals into executable experimental workflows.
@@
    LLM4GRN [Afonja2024LLM4GRN] integrates LLM reasoning with bioinformatics tools to discover causal gene regulatory networks, automating complex data analysis for biological mechanism inference.
@@
    Hierarchically Encapsulated Representation [Shi2025HierEncapRepProtocol] employs hierarchical architectures for multi-level protocol reasoning—from macro workflows to micro parameters—enabling robust, context-aware self-driving lab planning.

### v5

@@ 3.2 Idea Generation
Idea generation transforms structured knowledge into testable hypotheses through scientific creativity, semantic grounding, and novelty control. We formalize this as a four-stage framework (Figure 5): conceptual fusion, hypothesis refinement, multi-agent brainstorming, and feasibility evaluation [70, 53].
@@
    Stage 1: Conceptual Fusion and Trend Extrapolation. Discovers novel research directions by identifying latent connections across disparate knowledge domains. Dense retrieval surfaces conceptually related papers, embedding-based clustering reveals thematic patterns, and knowledge graph reasoning [70] enables analogical transfer. Temporal citation analysis identifies emerging trends. This constructs novelty-annotated hypothesis banks with estimated originality scores.
@@
    Stage 2: Hypothesis Refinement and Knowledge-Grounded Pruning. Filters and refines candidate hypotheses through multi-criteria validation ensuring logical coherence and feasibility. Knowledge graph validation checks consistency with theoretical frameworks, provenance-aware distillation [59] traces components to supporting evidence, and simulation-based testing evaluates tractability. Advanced systems employ reinforcement learning for iterative refinement and uncertainty-aware rejection sampling to eliminate speculative hypotheses.
@@
    Stage 3: Multi-Agent Brainstorming and Cross-Domain Recombination. Enhances hypothesis diversity through structured collaboration among specialized agents. Multi-role architectures assign distinct responsibilities—critic agents identify flaws, generator agents propose variations, verifier agents check consistency [53, 14]. Structured message passing enables agents to ground discussions in retrieved evidence while exploring complementary perspectives, producing hypothesis sets with conceptual diversity.
@@
    Stage 4: Feasibility and Novelty Evaluation. Ranks hypotheses through multi-dimensional scoring combining technical feasibility, scientific novelty, and potential impact. Graph-based semantic evaluation quantifies conceptual distance from existing work, while expert-annotated benchmarks [19, 33] provide ground-truth assessments. High-scoring hypotheses receive priority routing to experimental preparation, establishing a quality-controlled pipeline from ideation to experimentation.
@@ 3.3 Experimental Preparation
    Stage 1: Experimental Framing. Translates hypotheses into actionable experimental plans by formalizing variables, metrics, and testable assumptions. Multimodal reasoning over tables and charts [62] enables structured variable definition, while automated model template generation [39] provides standardized experimental frameworks. This stage establishes clear objectives, measurable outcomes, and explicit constraints for subsequent execution.
@@
    Stage 3: Protocol Implementation and Plan Reasoning. Operationalizes experiments through mathematical modeling and multi-step execution planning. Intermediate checkpoints [20] enable continuous introspection and error recovery, while dynamic schema adaptation allows real-time protocol adjustments. This stage transforms high-level experimental designs into executable sequences with built-in monitoring and adaptive re-execution capabilities.
@@
    Stage 4: Reproducibility and Lifecycle Tracking. Maintains comprehensive provenance documentation throughout experimental lifecycles. Parameter tracking, environmental logging [17], and versioning systems record all experimental configurations and execution details. This stage quantifies cross-agent consistency metrics and ensures verifiable scientific operations through cryptographic hashing, version control, and standardized metadata schemas.
@@ 3.4 Experimental Execution
Experimental execution operationalizes hypothesis-driven protocols into verifiable empirical outcomes through closed-loop reasoning and adaptive control [5, 24, 25, 40]. Unlike linear automation pipelines, this stage employs continuous feedback between symbolic planning and physical instrumentation, enabling real-time error correction and parameter optimization. We formalize this process through four sequential stages (Figure 7): Protocol Instantiation, Instrument Invocation, Adaptive Execution, and Data Validation.
@@
    Stage 1: Protocol Instantiation. Translates abstract experimental designs into executable, domain-specific procedures through language-to-action mapping. BioPlanner demonstrates natural-language-to-protocol conversion for biochemical validation [42], while Curie formalizes modular representations enabling task decomposition [24]. Hierarchical frameworks encode safety constraints, device mappings, and procedural templates into structured abstraction layers [50].
@@
    Stage 2: Instrument and Tool Invocation. Interfaces instantiated protocols with heterogeneous experimental infrastructure including robotic laboratories, simulators, and computational clusters. Self-driving laboratories [57] integrate robotic manipulation and closed-loop control, while multi-agent systems like ORGANA [9] and AutoLabs [43] orchestrate apparatus scheduling. Autonomous agents demonstrate scalability through unified API integration and metadata-driven control.
@@
    Stage 3: Adaptive Execution and Feedback. Implements real-time monitoring and dynamic protocol revision through closed-loop control. EXP-Bench [25] and Curie [24] leverage intermediate metrics to trigger parameter adjustments. DS-1000 [26] and notebook generation [67] demonstrate code-testing feedback loops, while CoScientist [5] realizes physical closed-loop reasoning. These frameworks converge reinforcement learning, self-critique, and error detection into adaptive control.
@@
    Stage 4: Data Acquisition and Validation. Consolidates multimodal experimental outputs into provenance-linked, validated datasets through comprehensive metadata annotation and statistical verification [40]. Curie [24] and EXP-Bench [25] embed runtime analytics validating outputs against simulations. Physics applications aggregate sensor telemetry for causal interpretation. This establishes an auditable evidence chain linking execution logs to scientific reporting.
@@ 3.5 Scientific Writing
Scientific writing transforms experimental findings into verifiable, publication-grade scholarly narratives through factual grounding, data–text alignment, and ethical compliance. LLM-assisted writing now permeates biomedical publishing, educational research [30], and simulation science [8]. We formalize this ecosystem through five interlinked stages (Figure 8): Drafting, Data–Text Linking, Peer-Review Automation, Ethical Governance, and Publication Optimization.
@@
    Stage 1: Drafting and Structural Composition. Converts analytical outputs into disciplinary-compliant academic prose through logical structuring and cross-sectional coherence [56, 23]. Domain-specific assistants achieve expert-level fluency via section-specific prompts [8], while citation-aware mechanisms map claims to verified sources and structure-aware templates enforce parallelism. Human-AI co-editing workflows reduce grammatical noise and accelerate drafting [18].
@@
    Stage 2: Data–Text Linking and Multimodal Representation. Aligns quantitative data with narratives through automated cross-referencing of tables and figures [63]. Hybrid language-vision models generate visual elements from structured data [47], while grounded captioning mechanisms automate caption rewriting and figure suggestion [51]. This establishes scientific writing as multimodal composition where LLMs mediate between data, visuals, and text.
@@
    Stage 3: Peer-Review Automation and Quality Verification. Automates manuscript evaluation through LLM-based reviewers trained on editorial corpora, identifying inconsistent claims and logical fallacies [69, 51]. Systems emulate referee workflows via structured comments and quantitative scores, implementing iterative review loops that reduce latency [12]. Reproducibility and algorithmic bias remain challenges, necessitating transparent confidence calibration [56].
@@
    Stage 4: Ethical Compliance and Authorship Governance. Enforces mandatory disclosure of AI involvement to prevent ghostwriting [4, 45, 41]. Standardized frameworks document prompt histories, model identifiers, and revision provenance [46], while layered attribution models recognize human intellectual contribution and algorithmic assistance [22]. These mechanisms anchor trust and accountability in AI-augmented authorship.
@@
    Stage 5: End-to-End Publication Optimization. Integrates drafting, multimodal generation, review, and ethics into unified workflows. Frameworks like WritingBench [63] and SPOT [51] evaluate manuscripts for factual accuracy and citation soundness. Adaptive agents monitor document lifecycles, suggesting figure adjustments, formatting references, and validating ethical disclosures, enabling scientists to transition to supervisory orchestration [23].
@@ 3.6 Paper Generation
Paper generation integrates ideation, experimentation, visualization, composition, and review into autonomous manuscript creation workflows (Figure 9). Systems like The AI Scientist v1/v2 generate complete papers—including figures, experiments, and reviews—with minimal human oversight [36, 66], while AI-Researcher establishes end-to-end pipelines and benchmarks (Scientist-Bench) for systematic evaluation [55]. We decompose this process into four stages: Manuscript Drafting, Visual & Tabular Composition, Review & Revision, and Publication Dissemination.
@@
    Stage 1: Manuscript Drafting. Transforms structured outputs into full manuscript sections through retrieval-augmented generation. The AI Scientist v1 converts experiment logs to LaTeX drafts, synthesizes literature context, and embeds citations automatically [36]. AIGS extends this via multi-agent pipelines for full-cycle generation [34], employing template-based formatting for journal compliance.
@@
    Stage 2: Visual & Tabular Composition. Generates and integrates visual artifacts with narrative alignment. The AI Scientist converts experimental outputs to plots and tables, generating captions and inserting cross-references [36], while SciSciGPT integrates data processing with figure generation [49]. Core technologies include automatic plotting, figure-caption co-generation, and multimodal consistency validation.
@@
    Stage 3: Review & Revision Agent. Embeds internal reviewer agents to critique and refine manuscripts. The AI Scientist v2 implements agentic tree-search with dedicated reviewer components generating structured feedback [66], while AI-Researcher enables automated revision loops [55]. Key methodologies include peer-review-trained models, LLM-driven comment generation, and iterative author-reviewer simulation.
@@
    Stage 4: Publication Dissemination. Prepares manuscripts for submission through venue-specific formatting and metadata management. Automated pipelines ensure submission readiness [38], while post-publication modules enable versioning and citation tracking. Core technologies include template-mapping tools, submission-API integration, metadata provenance logging, and ethical-disclosure verification.
@@ 4.1 General AI Scientist Systems
    The AI Scientist v1 [36] deploys modular multi-agent architecture—planning, coding, analysis, writing—coordinated by a meta-scientist module, autonomously selecting topics, generating executable code, drafting papers, and reproducing canonical machine learning studies.
@@
    The AI Scientist v2 [66] advances v1 through agentic tree-search planning, dynamically exploring parallel hypotheses and evaluating novelty/validity via reflective feedback loops.
@@
    AI-Researcher [55] prioritizes transparency through provenance-tracking memory graphs recording all artifacts—code, data logs—co-developed with Scientist-Bench for reproducibility evaluation.
@@
    Curie [24] achieves rigorous experimental control via causality-aware planning loops, automating machine learning hypothesis testing with explicit causal assumptions.
@@ 4.2 Chemistry and Materials Science
    Coscientist [5] integrates GPT-4 reasoning with robotic liquid handlers, autonomously planning reactions, generating control code, interpreting spectroscopic feedback, and iteratively refining hypotheses.
@@
    A-Lab [54] combines Bayesian optimization with LLM-guided experiment design, autonomously synthesizing and characterizing thousands of novel inorganic materials weekly.
@@
    Robotic AI Chemist [52] converges cognitive autonomy (literature-grounded reasoning) with physical autonomy (robotic manipulation) for end-to-end reaction design and execution.
@@
    AutoLabs [43] deploys multi-agent self-correction cycles—planning, control, safety auditing—detecting anomalies and recalibrating instruments for enhanced throughput and safety.
@@ 4.3 Biology and Biomedical Research
    BioPlanner [42] establishes benchmarks for LLM-driven biological protocol design, formalizing translation of research goals into executable experimental workflows.
@@
    LLM4GRN [1] integrates LLM reasoning with bioinformatics tools to discover causal gene regulatory networks, automating complex data analysis for biological mechanism inference.
@@
    Hierarchically Encapsulated Representation [50] employs hierarchical architectures for multi-level protocol reasoning—from macro workflows to micro parameters—enabling robust, context-aware self-driving lab planning.